OpenStudy (anonymous):

Find y if x is -3,-2,-1,0,1,2,3,4 Equation is y=2x

OpenStudy (anonymous):

it is easy just plug in the values of x let when x=-3 then y will be y=2(-3)=-6 now when x=-2 y=2(-2)=-4 i hope you can do this for other values of x .

OpenStudy (anonymous):

Okay but what if the equation is y=3x^2+6x+2

OpenStudy (anonymous):

same put the values of x for example if x=-3 then y=3(-3)^2+6(-2)+2 y=3(9)-12+2 y=27-12+2 y=17

OpenStudy (anonymous):

How did u get the -2